# From Cool Demos to Production-Ready FMware: Core Challenges and a Technology Roadmap

### v1

@@ IV-B Data and FM Alignment
Low data efficiency. Accurately estimating inbound and outbound data volume is crucial for production-ready FMware, impacting both budget and performance [42]. Underestimation leads to poor model performance, while overestimation wastes resources [43, 42].
IV-C Prompting
Prompting involves designing and refining prompts to guide FMs toward desired outputs using precise instructions and in-context learning. While demos can get away with simple or non-optimized prompts, production-ready FMware demands meticulous engineering for consistent, reliable performance [44].
God prompts. “God prompts” that try to handle multiple tasks at once lead to unpredictable outputs and complicate debugging and maintenance [21, 44]. In production, tasks should be split into smaller, focused prompts and be decoupled for better accuracy, debuggability and maintainability (similarly to code) [21].
Lack of Built-in QA Checks for Prompts. Production-ready FMware demands robust quality assurance (QA) for prompts, including semantic and structural validation. Without built-in QA checks, prompts may produce erroneous or inappropriate responses, leading to potential failures and a loss of user trust [45, 46, 47].
Lack of Prompt Portability Across FMs. Prompts optimized for one FM typically fail with others due to differing architectures and training data [48]. This drastically limits portability and poses challenges to software evolution.
Absence of FM-Specific Optimizations. FM-specific optimizations can boost performance by up to 10% by tailoring prompts to each model’s architecture [49]. However, these enhancements create dependencies, complicating updates and maintenance [50].
Complexity in determining failure rationale. Diagnosing whether failures result from prompt issues or FM shortcomings is difficult, often hindering the identification of the root cause and delaying effective resolution [51, 52].
Non-representative or insufficient in-context learning examples. Providing insufficient or non-representative examples for in-context learning degrades FM performance. For instance, Wang et al. [53] show that random in-context learning examples yield suboptimal outputs. Demos can use minimal examples, but production-ready FMware requires comprehensive, relevant examples for reliability.
IV-D Grounding
@@
Low information density in grounding data. Low-density grounding data impairs FM performance by limiting its ability to generate coherent responses [54, 55]. In production, rich, well-structured data is essential for accurate outputs. Laban et al. [54] show how poorly structured, vast information leads to failures in long-context tasks, while Wang et al. [55] emphasize efficient data compression and selection to maintain performance and relevance in complex queries.
Irrelevant grounding data. Incorporating irrelevant grounding data confuses FMs and reduces performance. Liu et al. [56] show that models struggle with long-context information, especially when crucial data is buried, while Cuconasu et al. [57] note how noisy data worsens the focus of retrieval systems. For production environments, filtering and curating data is crucial to ensure only pertinent information is used.
Overcomplicating solutions with advanced techniques. Using complex methods when simpler ones would suffice adds unnecessary overhead. For instance, neural retrieval can raise computational costs without added benefits when a keyword search would be enough [21, 58]. In production, simplicity promotes scalability and maintainability, while demos often prioritize novelty [21, 58].
IV-E Agent(s)
@@
God Agents. Monolithic “God agents” handling many tasks create complexity and maintenance challenges in production. In production, smaller specialized agents are preferred for modularity and reliability, offering better scalability and maintainability [59, 60].
@@
Too Low-Level Tool Usage Abstraction. Agents that require developers to interact with low-level tools increase system complexity and hinder scalability in production [61]. When agents demand manual management of API calls or intricate state transitions, they become error-prone and harder to maintain. High-level abstractions, which hide tool interactions, are crucial for reducing bugs and allowing developers to focus on business logic rather than technical details.
@@
Capability-Centric Instead of Use-Case-Centric Documentation for Tools. Documentation that focuses on tool capabilities rather than specific use cases makes it harder for FMs to leverage these tools. FMs need clear, practical examples to better operate, as seen in adaptive systems like Voyager, where the agents excel at tool use by leveraging concrete use cases. Without this, the learning curve steepens, increasing the risk of errors and slowing agent evolution [62, 61, 60]. In production, use-case-driven guidance is crucial for building reliable systems [62].
IV-F Agent Orchestration
@@
Orchestrating multiple agents in FMware, both FM-based and traditional, ensures efficient handling of complex tasks by assigning sub-tasks to specialized agents, improving scalability and reliability [59]. This mirrors microservices, where modular components collaborate for a common goal, integrating flexibility and maintainability into production systems.
@@
Over-reliance on FM planning capability. Relying solely on an FM’s planning capabilities in production FMware introduces risks like unpredictability and lack of transparency, often leading to errors and unreliable outcomes due to issues such as hallucinations or failure to align with business logic [63, 64]. In contrast, demos may succeed with FM-based planning due to their controlled environment.
IV-G Guarding
@@
Guarding ensures that FM and Agents using FM outputs are safe, accurate, and policy-compliant. Guardrails, for example, implements “rails” for input moderation, fact-checking, and hallucination control, maximizing reliability in production FMware [65, 66]. While demos may ignore these precautions, production systems must prioritize guardrails to ensure compliance, trust, and error prevention [65, 66].
@@
Simple keyword-based guarding is ineffective. Relying solely on keyword-based guarding is ineffective in production environments. Simple keyword filters can be easily bypassed or mistakenly block legitimate content due to language ambiguity. In production-ready FMware, sophisticated guardrails such as fact-checking, semantic filtering, or policy-guided decoding are essential for discerning intent and context, reducing both false positives and false negatives [65, 66].
@@
Lack of hallucination guardrails. Hallucinations, pose significant risks in production systems, especially in domains like healthcare or law, where factual accuracy is crucial. These fabricated outputs can lead to the spread of misinformation, undermine user trust, and even trigger legal liabilities [67, 68]. Additionally, FMs often lack confidence estimation, which makes it difficult to flag uncertain outputs for human review before they cause harm. User feedback systems, though essential for improving accuracy over time, are often underutilized, preventing continuous correction of inaccuracies [67, 68, 69].
IV-H FMware System Testing and Optimization
@@
Moreover, in production-ready FMware, system-wide testing must account for the dynamic interactions between agents, scalability concerns, and maintaining robustness under varying workloads. These challenges demand rigorous optimization, as the FMware needs to handle real-world complexities like fluctuating performance, inconsistent results, and unexpected edge cases. This goes beyond the controlled environments seen in demos, where simplicity and consistency are more achievable [70].
@@
Lack of latency handling mechanisms. Latency is a major challenge for production-ready FMware, leading to slow responses or timeouts under real-world conditions. Ahmed et al. [71] highlight how resource-heavy models like BERT struggle with responsiveness.  Santilli et al. [72] show that the sequential nature of model inference introduces delays, making real-time performance difficult [71, 72]. Variable loads further complicate the issue, as latency spikes during high demand can render systems unresponsive [73].
@@
Lack of Retry Optimizations. Failure to implement intelligent retry strategies can significantly reduce FMware reliability, particularly when interacting with cloud-hosted FMs. In production, transient errors are common, and retry storms—where repeated requests overwhelm a system—can be costly, leading to increased error rates and decreased availability [74, 75]. FMware also requires intelligent retries to handle scenarios like re-prompting for self-consistency, where multiple attempts may be needed to ensure the correct output [76].
@@
High Cost of Regression Testing. Regression testing in FMware is both expensive and time-consuming due to the non-deterministic nature of FMs and the complexity of their interactions. As FM APIs evolve, performance can degrade silently, making it essential to frequently re-evaluate across numerous scenarios to ensure system reliability. This increases costs and slows down development cycles. Furthermore, traditional regression testing methods—relying on exact output matching—are inadequate for FMware, which often produce variable results. More nuanced evaluation techniques, such as similarity and correctness testing, are required to detect subtle issues in behavior. This adds to the overall complexity and cost, especially when considering large-scale applications with significant feature sets [70, 77, 78]. In addition, the need for continuous evaluation across various system components in production environments demands more sophisticated testing setups, which are critical for maintaining system stability but increase both the time and resources required [70, 77].
@@
Absence of Software Performance Engineering Practices. The absence of established software performance engineering practices, such as model swapping (replacing on FM with another to optimize memory usage), significantly complicates the optimization and scaling of FMware. In production environments, model swapping requires careful planning and compatibility testing to ensure performance gains. Without these measures, FMware risks suboptimal resource utilization, leading to increased operational costs. For instance, Li et al. [79] and Huang et al. [80] stress the importance of system-level improvements in FM serving, including resource-efficient FM management to handle increasing demands. Additionally, Unknown [81] highlight how inadequate optimization strategies can bottleneck FM inference, further reinforcing the need for robust practices.
@@
Lack of controlled execution mechanisms. The lack of controlled execution mechanisms significantly complicates the verification of fixes and tracking the impact of updates in FMware. Without tools like feature flags, staging environments, or canary releases, testing changes before full deployment becomes risky, as updates can introduce new issues or break functionality [82]. FMware systems, like those discussed in SecGPT [82], emphasize the importance of isolating updates and managing risks, particularly in complex environments where multiple components interact.
IV-I Deployment and Maintenance
@@
Lack of Efficient Feedback Technology. Collecting and integrating feedback seamlessly is essential for refining models, addressing issues, and aligning with user needs. Without these mechanisms, developers miss critical insights into user behavior, leading to stagnation in performance and adaptability. Luo et al. [83] demonstrate the importance of a robust feedback loop in the presence of continuous FM and FMware evolution, showing how the absence of efficient data collection delays improvements by failing to identify and address FM weaknesses. Production environments, unlike demos, require automated, iterative feedback systems to enable continuous learning, performance optimization, and user engagement [84, 83, 85].
@@
Lack of FMware-native observability. The lack of FMware-native observability makes it difficult to diagnose issues, optimize performance, and ensure compliance in production-ready FMware [1]. Traditional observability tools focus primarily on functional observability metrics, such as latency, execution traces, and throughput, which are insufficient for capturing the internal reasoning and decision-making processes of FMs [1]. This opacity hinders the ability to identify root causes of failures, especially when errors stem from complex reasoning or coordination between agents. In addition, regulatory compliance often demands detailed logs and auditing, which are difficult to maintain without comprehensive observability.
@@
Ineffective FM Update Mechanisms Maintaining and updating FMs in production FMware presents significant issues compared to traditional software [1]. Fixes can take months, and updates are non-deterministic—providing more training data does not ensure specific issues are resolved. Improvements may not work as expected, and previously stable features can break. In addition, improvements noted in release notes do not directly translate to features, making it difficult for developers to test what has changed or verify the improved capability. This unpredictability complicates testing and maintenance, making it difficult to prioritize efforts. In production FMware, these issues undermine reliability, performance, and user trust.
IV-J Memory Management
@@
Memory management in FMware plays a crucial role across all stages of the lifecycle, including data alignment, prompting, grounding, agent orchestration, system optimization, and deployment. Effective memory management ensures that FMs access, store, and update relevant context and knowledge, impacting how models handle inference latency, retries, and scalability [86, 79, 81].
@@
During deployment, memory management supports model swapping, where efficient context handling prevents performance degradation from latency issues [81]. Additionally, memory is essential for managing retries by preserving context, preventing costly retry storms, and maintaining performance under load [75, 83]. It also ensures that agents share state and context, crucial for coherent decision-making across agents. Robust memory management is key for optimizing dynamic workloads and ensuring regulatory compliance [1, 84].
@@
Inefficient Knowledge Representation. Inefficient knowledge representation in FMware leads to storing redundant or irrelevant data, inflating computational costs and response times. For instance, Guo et al. [87] highlight that poor knowledge representation can cause inconsistent reasoning and incorrect outputs when retrieval systems are not optimized. In addition, the irrelevant information retrieved due to inefficient knowledge representation skew FM’s outputs [88].
@@
Cumbersome and Error-Prone In-Memory and Across-Memories Knowledge Management. Managing memory across multiple systems in FMware is complex and prone to errors such as synchronization issues and data corruption. Xie et al. [89] and Chen et al. [90] emphasize that conflicting knowledge within memory systems can lead to inconsistencies and errors when FMs must navigate between contradictory pieces of information, further complicating memory management in production environments. Packer et al. [91] highlight that production-ready FMware requires robust memory systems, as context loss or memory mismanagement can degrade user experiences or lead to system failures [91].
IV-K Cross-cutting Issues that affect productionizing FMware

### v2

@@ 4.2. Data and FM Alignment
Low data efficiency. Accurately estimating inbound and outbound data volume is crucial for production-ready FMware, impacting both budget and performance (Chen et al., 2023). Underestimation leads to poor model performance, while overestimation wastes resources (Wang et al., 2022; Chen et al., 2023).
4.3. Prompting
Prompting involves designing and refining prompts to guide FMs toward desired outputs using precise instructions and in-context learning. While demos can get away with simple or non-optimized prompts, production-ready FMware demands meticulous engineering for consistent, reliable performance (Parthasarathy et al., 2024).
God prompts. “God prompts” that try to handle multiple tasks at once lead to unpredictable outputs and complicate debugging and maintenance (Yan et al., [n. d.]; Parthasarathy et al., 2024). In production, tasks should be split into smaller, focused prompts and be decoupled for better accuracy, debuggability and maintainability (similarly to code) (Yan et al., [n. d.]).
Lack of Built-in QA Checks for Prompts. Production-ready FMware demands robust quality assurance (QA) for prompts, including semantic and structural validation. Without built-in QA checks, prompts may produce erroneous or inappropriate responses, leading to potential failures and a loss of user trust (Kamoi et al., 2024; Ahmed and Devanbu, 2023; Wang et al., 2024b).
Lack of Prompt Portability Across FMs. Prompts optimized for one FM typically fail with others due to differing architectures and training data (Chang et al., 2024). This drastically limits portability and poses challenges to software evolution.
Absence of FM-Specific Optimizations. FM-specific optimizations can boost performance by up to 10% by tailoring prompts to each model’s architecture (Chen et al., 2024c). However, these enhancements create dependencies, complicating updates and maintenance (Sabbatella et al., 2024).
Complexity in determining failure rationale. Diagnosing whether failures result from prompt issues or FM shortcomings is difficult, often hindering the identification of the root cause and delaying effective resolution (Ashtari et al., 2023; Amujo and Yang, 2024).
Non-representative or insufficient in-context learning examples. Providing insufficient or non-representative examples for in-context learning degrades FM performance. For instance, Wang et al. (2023c) show that random in-context learning examples yield suboptimal outputs. Demos can use minimal examples, but production-ready FMware requires comprehensive, relevant examples for reliability.
4.4. Grounding
@@
Low information density in grounding data. Low-density grounding data impairs FM performance by limiting its ability to generate coherent responses (Laban et al., 2024; Wang et al., 2024c). In production, rich, well-structured data is essential for accurate outputs. Laban et al. (2024) show how poorly structured, vast information leads to failures in long-context tasks, while Wang et al. (2024c) emphasize efficient data compression and selection to maintain performance and relevance in complex queries.
Irrelevant grounding data. Incorporating irrelevant grounding data confuses FMs and reduces performance. Liu et al. (2024a) show that models struggle with long-context information, especially when crucial data is buried, while Cuconasu et al. (2024) note how noisy data worsens the focus of retrieval systems. For production environments, filtering and curating data is crucial to ensure only pertinent information is used.
Overcomplicating solutions with advanced techniques. Using complex methods when simpler ones would suffice adds unnecessary overhead. For instance, neural retrieval can raise computational costs without added benefits when a keyword search would be enough (Yan et al., [n. d.]; Sawarkar et al., 2024). In production, simplicity promotes scalability and maintainability, while demos often prioritize novelty (Yan et al., [n. d.]; Sawarkar et al., 2024).
4.5. Agent(s)
@@
God Agents. Monolithic “God agents” handling many tasks create complexity and maintenance challenges in production. In production, smaller specialized agents are preferred for modularity and reliability, offering better scalability and maintainability (Rasal and Hauer, 2024; Chen et al., 2024b).
@@
Too Low-Level Tool Usage Abstraction. Agents that require developers to interact with low-level tools increase system complexity and hinder scalability in production (Yang et al., 2024). When agents demand manual management of API calls or intricate state transitions, they become error-prone and harder to maintain. High-level abstractions, which hide tool interactions, are crucial for reducing bugs and allowing developers to focus on business logic rather than technical details.
@@
Capability-Centric Instead of Use-Case-Centric Documentation for Tools. Documentation that focuses on tool capabilities rather than specific use cases makes it harder for FMs to leverage these tools. FMs need clear, practical examples to better operate, as seen in adaptive systems like Voyager, where the agents excel at tool use by leveraging concrete use cases. Without this, the learning curve steepens, increasing the risk of errors and slowing agent evolution (Wang et al., 2023b; Yang et al., 2024; Chen et al., 2024b). In production, use-case-driven guidance is crucial for building reliable systems (Wang et al., 2023b).
4.6. Agent Orchestration
@@
Orchestrating multiple agents in FMware, both FM-based and traditional, ensures efficient handling of complex tasks by assigning sub-tasks to specialized agents, improving scalability and reliability (Rasal and Hauer, 2024). This mirrors microservices, where modular components collaborate for a common goal, integrating flexibility and maintainability into production systems.
@@
Over-reliance on FM planning capability. Relying solely on an FM’s planning capabilities in production FMware introduces risks like unpredictability and lack of transparency, often leading to errors and unreliable outcomes due to issues such as hallucinations or failure to align with business logic (Guo et al., 2024a; Huang et al., 2024a). In contrast, demos may succeed with FM-based planning due to their controlled environment.
4.7. Guarding
@@
Guarding ensures that FM and Agents using FM outputs are safe, accurate, and policy-compliant. Guardrails, for example, implements “rails” for input moderation, fact-checking, and hallucination control, maximizing reliability in production FMware (Rebedea et al., 2023; Dong et al., 2024). While demos may ignore these precautions, production systems must prioritize guardrails to ensure compliance, trust, and error prevention (Rebedea et al., 2023; Dong et al., 2024).
@@
Simple keyword-based guarding is ineffective. Relying solely on keyword-based guarding is ineffective in production environments. Simple keyword filters can be easily bypassed or mistakenly block legitimate content due to language ambiguity. In production-ready FMware, sophisticated guardrails such as fact-checking, semantic filtering, or policy-guided decoding are essential for discerning intent and context, reducing both false positives and false negatives (Rebedea et al., 2023; Dong et al., 2024).
@@
Lack of hallucination guardrails. Hallucinations, pose significant risks in production systems, especially in domains like healthcare or law, where factual accuracy is crucial. These fabricated outputs can lead to the spread of misinformation, undermine user trust, and even trigger legal liabilities (Ayyamperumal and Ge, 2024; Dahl et al., 2024). Additionally, FMs often lack confidence estimation, which makes it difficult to flag uncertain outputs for human review before they cause harm. User feedback systems, though essential for improving accuracy over time, are often underutilized, preventing continuous correction of inaccuracies (Ayyamperumal and Ge, 2024; Dahl et al., 2024; Spiess et al., 2024).
4.8. FMware System Testing and Optimization
@@
Moreover, in production-ready FMware, system-wide testing must account for the dynamic interactions between agents, scalability concerns, and maintaining robustness under varying workloads. These challenges demand rigorous optimization, as the FMware needs to handle real-world complexities like fluctuating performance, inconsistent results, and unexpected edge cases. This goes beyond the controlled environments seen in demos, where simplicity and consistency are more achievable (Ma et al., 2024).
@@
Lack of latency handling mechanisms. Latency is a major challenge for production-ready FMware, leading to slow responses or timeouts under real-world conditions. Ahmed et al. (2022) highlight how resource-heavy models like BERT struggle with responsiveness.  Santilli et al. (2023) show that the sequential nature of model inference introduces delays, making real-time performance difficult (Ahmed et al., 2022; Santilli et al., 2023). Variable loads further complicate the issue, as latency spikes during high demand can render systems unresponsive (Stamatescu, [n. d.]).
@@
Lack of Retry Optimizations. Failure to implement intelligent retry strategies can significantly reduce FMware reliability, particularly when interacting with cloud-hosted FMs. In production, transient errors are common, and retry storms—where repeated requests overwhelm a system—can be costly, leading to increased error rates and decreased availability (sca, [n. d.]; RobBagby, [n. d.]). FMware also requires intelligent retries to handle scenarios like re-prompting for self-consistency, where multiple attempts may be needed to ensure the correct output (sig, [n. d.]).
@@
High Cost of Regression Testing. Regression testing in FMware is both expensive and time-consuming due to the non-deterministic nature of FMs and the complexity of their interactions. As FM APIs evolve, performance can degrade silently, making it essential to frequently re-evaluate across numerous scenarios to ensure system reliability. This increases costs and slows down development cycles. Furthermore, traditional regression testing methods—relying on exact output matching—are inadequate for FMware, which often produce variable results. More nuanced evaluation techniques, such as similarity and correctness testing, are required to detect subtle issues in behavior. This adds to the overall complexity and cost, especially when considering large-scale applications with significant feature sets (Ma et al., 2024; sco, 2023; hum, 2023). In addition, the need for continuous evaluation across various system components in production environments demands more sophisticated testing setups, which are critical for maintaining system stability but increase both the time and resources required (Ma et al., 2024; sco, 2023).
@@
Absence of Software Performance Engineering Practices. The absence of established software performance engineering practices, such as model swapping (replacing on FM with another to optimize memory usage), significantly complicates the optimization and scaling of FMware. In production environments, model swapping requires careful planning and compatibility testing to ensure performance gains. Without these measures, FMware risks suboptimal resource utilization, leading to increased operational costs. For instance, Li et al. (2024) and Huang et al. (2024b) stress the importance of system-level improvements in FM serving, including resource-efficient FM management to handle increasing demands. Additionally, Łazuka et al. (2024) highlight how inadequate optimization strategies can bottleneck FM inference, further reinforcing the need for robust practices.
@@
Lack of controlled execution mechanisms. The lack of controlled execution mechanisms significantly complicates the verification of fixes and tracking the impact of updates in FMware. Without tools like feature flags, staging environments, or canary releases, testing changes before full deployment becomes risky, as updates can introduce new issues or break functionality (Wu et al., 2024a). FMware systems, like those discussed in SecGPT (Wu et al., 2024a), emphasize the importance of isolating updates and managing risks, particularly in complex environments where multiple components interact.
4.9. Deployment and Maintenance
@@
Lack of Efficient Feedback Technology. Collecting and integrating feedback seamlessly is essential for refining models, addressing issues, and aligning with user needs. Without these mechanisms, developers miss critical insights into user behavior, leading to stagnation in performance and adaptability. Luo et al. (2024) demonstrate the importance of a robust feedback loop in the presence of continuous FM and FMware evolution, showing how the absence of efficient data collection delays improvements by failing to identify and address FM weaknesses. Production environments, unlike demos, require automated, iterative feedback systems to enable continuous learning, performance optimization, and user engagement (Wang et al., 2023a; Luo et al., 2024; aws, [n. d.]).
@@
Lack of FMware-native observability. The lack of FMware-native observability makes it difficult to diagnose issues, optimize performance, and ensure compliance in production-ready FMware (Hassan et al., 2024b). Traditional observability tools focus primarily on functional observability metrics, such as latency, execution traces, and throughput, which are insufficient for capturing the internal reasoning and decision-making processes of FMs (Hassan et al., 2024b). This opacity hinders the ability to identify root causes of failures, especially when errors stem from complex reasoning or coordination between agents. In addition, regulatory compliance often demands detailed logs and auditing, which are difficult to maintain without comprehensive observability.
@@
Ineffective FM Update Mechanisms Maintaining and updating FMs in production FMware presents significant issues compared to traditional software (Hassan et al., 2024b). Fixes can take months, and updates are non-deterministic—providing more training data does not ensure specific issues are resolved. Improvements may not work as expected, and previously stable features can break. In addition, improvements noted in release notes do not directly translate to features, making it difficult for developers to test what has changed or verify the improved capability. This unpredictability complicates testing and maintenance, making it difficult to prioritize efforts. In production FMware, these issues undermine reliability, performance, and user trust.
4.10. Memory Management
@@
Memory management in FMware plays a crucial role across all stages of the lifecycle, including data alignment, prompting, grounding, agent orchestration, system optimization, and deployment. Effective memory management ensures that FMs access, store, and update relevant context and knowledge, impacting how models handle inference latency, retries, and scalability (Kwon et al., 2023; Li et al., 2024; Łazuka et al., 2024).
@@
During deployment, memory management supports model swapping, where efficient context handling prevents performance degradation from latency issues (Łazuka et al., 2024). Additionally, memory is essential for managing retries by preserving context, preventing costly retry storms, and maintaining performance under load (RobBagby, [n. d.]; Luo et al., 2024). It also ensures that agents share state and context, crucial for coherent decision-making across agents. Robust memory management is key for optimizing dynamic workloads and ensuring regulatory compliance (Hassan et al., 2024b; Wang et al., 2023a).
@@
Inefficient Knowledge Representation. Inefficient knowledge representation in FMware leads to storing redundant or irrelevant data, inflating computational costs and response times. For instance, Guo et al. (2024b) highlight that poor knowledge representation can cause inconsistent reasoning and incorrect outputs when retrieval systems are not optimized. In addition, the irrelevant information retrieved due to inefficient knowledge representation skew FM’s outputs (Wu et al., 2024b).
@@
Cumbersome and Error-Prone In-Memory and Across-Memories Knowledge Management. Managing memory across multiple systems in FMware is complex and prone to errors such as synchronization issues and data corruption. Xie et al. (2023) and Chen et al. (2022) emphasize that conflicting knowledge within memory systems can lead to inconsistencies and errors when FMs must navigate between contradictory pieces of information, further complicating memory management in production environments. Packer et al. (2023) highlight that production-ready FMware requires robust memory systems, as context loss or memory mismanagement can degrade user experiences or lead to system failures (Packer et al., 2023).
4.11. Cross-cutting Issues that affect productionizing FMware